OpenStudy (anonymous):

step by step help please 4m − 8 = 2m

OpenStudy (uri):

4m-8=2m 4m-2m=8 2m=8 m=8/2 m=4

OpenStudy (mathstudent55):

You need to combine all terms with m together first. To do that, first subtract 2m from both sides: 4m - 8 = 2m -2m -2m -------------- 2m - 8 = 0 Now you add 8 to both sides: 2m - 8 = 0 +8 +8 -------------- 2m = 8 Now divide both sides by 2 to get m by itself: \( \dfrac{2m}{2} = \dfrac{8}{2} \) \(m = 4\) Solution: \(m = 4\)